OpenStudy (anonymous):

Hey guys help me out with this and it'll be vastly appreciated. :) Giving medals.

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

\[y = \sqrt{4-x} + 3\] explain if its a function find the domain find the range

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well domain of a function is all the numbers that can be used in the function.perhaps in this function we can't find f(5) .

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

so 4-x must be larger or equal to0: \[4 - x \ge 0\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@SeverelyAutistic

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

well,did you get it?

OpenStudy (anonymous):

\[4 - x \ge 0\]

OpenStudy (anonymous):

Yea, is the vertical line test needed at all here?

OpenStudy (anonymous):

so \[x \le 4\]

OpenStudy (anonymous):

it's the domain ;)

OpenStudy (anonymous):

How would I know if its a function?

OpenStudy (anonymous):

And yes thank you for the domain part I understood perfectly :D

OpenStudy (anonymous):

you know this: 

OpenStudy (anonymous):

Don't think I do... could you explain?

OpenStudy (anonymous):

you're welcome ;) here if for two x we get the same f(x) it IS NOT a function.

OpenStudy (anonymous):

Ahhh So would I need to plug in values and check to see if they have the same results?

OpenStudy (anonymous):

no it's not necessary.you can find if it's a function or not easily. for example sqr (4-x) is cannot give us the same f(x) despite the Xs are different!

OpenStudy (anonymous):

you get what am i saying?

OpenStudy (anonymous):

Yep :D So we got the domain, how about the range?

OpenStudy (anonymous):

well,range is all the numbers the function give us! sqr (4-x) can give us at least 0 ... absolutely we cannot get -2 or -4 from sqr (4-x) !

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Hm, so what you're saying is that sqr(4-X), you will always get a positive value, and thus it wouldn't satisfy the domain?

OpenStudy (anonymous):

Took me a minute but I think I got it

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes ;) sry i want to go...so the sqr (4-x) only give us 0,1,2 and etc...so we'll get at least 0 and we have 3 too, so we'll get at least 3 from the function...so the range is: \[f(x) \ge 3\]

OpenStudy (aum):

Or, in interval notation, the range is: [3, infinity)

OpenStudy (anonymous):

Alright thank you so much, have a medal :)

OpenStudy (anonymous):

you're welcome ;)

OpenStudy (anonymous):

thanks for medal :)